OpenStudy (anonymous):

solve for x (x + 7)(x − 5)/2 = 32

OpenStudy (anonymous):

In order to solve for x, you need to find the roots, so you need to make the right side zero. (x+7)(x-5)=64 x^2+7x-5x-35=64 x^2+2x-99=0 After making it zero, you can just factor the equation, to get the answer. (x-9)(x+11)=0 x=9,-11